OpenStudy (anonymous):

subtract the fraction and reduce (x/(x^2-5x+6)) - (2/(x^2-4))

OpenStudy (anonymous):

Can you factorize this one: x^2-5x+6 = ...?

OpenStudy (anonymous):

(x-3) (x-2) and (x+2)(x-2)

OpenStudy (anonymous):

and i know you need like terms

OpenStudy (anonymous):

Great, so now you will have the common denominator is ( x+3) (x+2) ( x-2) Can you continue?

OpenStudy (anonymous):

Actually x^2-5x+6 = (x-3) (x+ 2)

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

i got an answer but apparently it was wrong i was hoping you could work it out so i could see where i went wrong.... my answer was

OpenStudy (anonymous):

Oop, hold on!

OpenStudy (anonymous):

Since common denominator is ( x-3) (x+2) ( x-2): x (x-2) = x² - 2x 2 ( x-3) = 2x - 6 Subtract them: -> (x² - 4x+ 6 ) / ( x-3) (x+2) ( x-2):

OpenStudy (anonymous):

and you can't factor the to[ right

OpenStudy (anonymous):

top****

OpenStudy (anonymous):

That's final result!

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

So do you know where you're confused?

OpenStudy (anonymous):

yea i tryed to factor out the top and you can't do that thats why i go tit wrong

OpenStudy (anonymous):

Because the top has Δ < 0, and the question doesn't mention about imaginary solution!